OpenStudy (anonymous):

integrate from 0 to 8. sqroot of (1+4/9x^-2/3) dx

OpenStudy (anonymous):

\[\sqrt{1+ \frac{4}{9} x^{-\frac{2}{3} } }\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

factor out the (4/9) , then do a trig substitution

OpenStudy (anonymous):

\[= \sqrt{ \frac{4}{9} ( \frac{9}{4} + x^{-\frac{2}{3} } )} = \frac{2}{3} \sqrt{ \frac{9}{4} +x^{-\frac{2}{3}}}\]

OpenStudy (anonymous):

now sub x= (9/4) tan(theta) or something similiar.

OpenStudy (anonymous):

or maybe tan^3

OpenStudy (anonymous):

uii